) What is the equilibrium price and quantity of fertilizer in an unregulated, competitive market? (4 marks) b) What is the efficient quantity of fertilizer? (4 marks) c) Suppose government imposes a tax equal to the marginal external cost. What is the equilibrium price paid by consumers and the equilibrium quantity after implementation of the tax? (6 marks) d) At the output level in part (c), how much is the tax? (5 marks) e) How much tax revenue does government collect? (6 marks) f) What is the deadweight loss borne by society if the externality is left uncorrected?


1
Expert's answer
2021-04-28T20:13:17-0400

a)Equilibrium price of fertilizer in an unregulated competitive market occurs when the price is at "1000" while quantity is at 6,that is where marginal cost and marginal benefit intersect.

b)The efficient quantity of fertilizer is found when; marginal social cost"=" marginal benefit, that is 4 tons per day.

c)Equilibrium price paid by consumers after implementation of the tax is "1200" and the equilibrium quantity after implementation of the tax is 4.

d)The tax at the output level in part (c)is; "1200-800=400".

e)The tax revenue the government collects is; tax"*" quantity

"4*400=1600".

f)The deadweight loss borne by society if the externality is left uncorrected is;"{1\over2}*(400)*(6-4)=400".